# CheMFi: A Multifidelity Dataset of Quantum Chemical Properties of Diverse Molecules

### v1

@@ 4.1 Single Molecule Benchmarks
A similar benchmarking procedure was carried out for the prediction of first vertical excitation energies of o-HBDI. In this case, the unsorted CM were used with the Matérn Kernel. Further details are presented in Section S2 of the supplementary file. As for the case of SMA, a preliminary analysis study was performed with the resulting plots shown in Figure 3(a). The difference in fidelities plot in the center indicates that the assumed hierarchy holds true for the fidelities. However, the fidelity scatter plot on the right hand side shows two distinct clusters. These correspond to the two main conformers of o-HBDI, namely the cis and trans conformers. The scatter also shows some cases where the STO3G fidelity covers a wider range of values and is less localized than the other fidelities. This could indicate that the use of STO3G in the MFML models would result in a lower improvement of the accuracy of the model.
The learning curves for the prediction of the first vertical excitation energies of o-HBDI from CheMFi are shown in Figure 3(b). The MAE are reported in $\rm cm^{-1}$ with the axes identically scaled for both MFML and o-MFML. With the addition of cheaper fidelities, the learning curves show a constant reduction in the offset of the MAE as seen in the near parallel learning curves of the different baselines fidelities. As anticipated from the preliminary analysis, the addition of the STO3G fidelity does not provide significant improvement especially for larger training samples. However, this is rectified, as expected, by the o-MFML method which was indeed shown to fix this very issue in Ref. 15.
4.2 Cumulative use of the dataset
Figure 4: Learning curves for MFML and o-MFML for SCF energies based on the compounded use of the CheMFi dataset. 1,500 samples were randomly chosen from each molecule to perform this example test.
The CheMFi dataset contains multifidelity QC properties of 9 molecules for 15,000 geometries. This totals to $9\times 15,000=135,000$ point calculations of the QC properties. This is therefore the largest collection of multifidelity dataset which can be used in various benchmarking processes. To demonstrate this form of cumulative use of the dataset, multifidelity models from Ref. 15 were tested against this in predicting the SCF of the molecules. From each molecule of the CheMFi dataset, 1,500 geometries were randomly chosen and compiled into a total of $9\times 1,500=13,500$ data points. From the 13,500 samples, a random set of 11,000 samples were used as the multifidelity training data. Of the remaining, 500 samples were used as a validation set and 2,000 as the holdout test set. With this setup learning curves were generated for the different multifidelity models in the same fashion as prescribed in Ref. 15.
The results of this test are shown in Figure 4 for MFML and o-MFML models. The learning curves show a decreasing slope for both cases for the different baseline fidelities. The addition of each cheaper fidelity results in a lower offset of MAE. The constant slope on the log-log axis indicates that addition of training samples can further decrease the MAE. On the right-hand side plot the scatter of reference TZVP versus MFML predicted SCF values are delineated. Across the energy ranges the MFML model predicts the SCF energies accurately as can be inferred from the scatter of the values being close to the identity mapping line.
5 Usage Notes
In addition to the multifidelity dataset, various tools to assess and benchmark multifidelity methods are also provided. These include scripts to perform preliminary analysis of the data based on the property of choice as recommended in Ref. 13, and the scripts to produce learning curves. Further, scripts to generate unsorted Coulomb Matrices, and the global SLATM descriptors are provided which are built upon the qmlcode package 29. The scripts are easy to use and well documented allowing for a streamlined benchmarking process with an example shown in Listing LABEL:lst_prelimanalysis for the preliminary analysis. Listing LABEL:lst_MFML_LCs shows an example to generate the multifidelity learning curves.
@@ 5 Usage Notes
Listing 2: Python example to perform preliminary analysis of SCF multifidelity data for the SMA molecule.
@@
Listing 3: Python example to generate multifidelity learning curves and the corrsponding plot of SCF for the SMA molecule.
Associated Content
Supplementary sections S1-S2, Figure S1.
Data availability
The dataset is hosted as an open sourced and citable repository on Zenodo at the URL https://zenodo.org/records/11636903.
Code availability
Acknowledgments
The authors acknowledge support by the DFG through projects ZA 1175/3-1 as well as through the DFG Priority Program SPP 2363 on “Utilization and Development of Machine Learning for Molecular Applications – Molecular Machine Learning”. The computations of the quantum chemical properties were carried out on the PLEIADES cluster at the University of Wuppertal, which was supported by the Deutsche Forschungsgemeinschaft (DFG, grant No. INST 218/78-1 FUGG) and the Bundesministerium für Bildung und Forschung (BMBF). The authors would also like to thank the ‘Interdisciplinary Center for Machine Learning and Data Analytics (IZMD)’ at the University of Wuppertal.
Declarations

### v2

@@ Acknowledgements
Figure 5: Learning curves for MFML and o-MFML for SCF ground state energies based on the cumulative use of the CheMFi dataset. 1,500 samples were randomly chosen from each molecule to perform this example test. The single fidelity KRR is also shown for comparison. The scatter plot of MFML predicted versus reference ground state energies is also shown. The MFML and o-MFML models perform well on the cumulative dataset showing errors in the range of a few hE for the ground state energies.
Figure 6: MAE versus time taken to generate a training set for MFML and o-MFML for the cumulative use of the CheMFi dataset. Here one observes the time-benefit of using MFML over single fidelity KRR for each cheaper baseline fidelity.
@@
Listing 2: Python example to perform preliminary analysis of SCF ground state multifidelity data for the SMA molecule.
@@
Listing 3: Python example to generate multifidelity learning curves and the corrsponding plot of SCF ground state energies for the SMA molecule.